In Yvette's bank, there is $2.60 in pennies, nickels, and dimes. IN ALL, THERE are 45 coins. If there are twice as many nickles as pennies, how many coins of each kind are there in the bank?
Answer by stanbon(48551)

In Yvette's bank, there is $2.60 in pennies, nickels, and dimes. IN ALL, THERE are 45 coins. If there are twice as many nickles as pennies, how many coins of each kind are there in the bank?
-----------------
Let number of pennies be "x"; Value of these is x cents
Number of nickels "2x" ; Value of these is 5(2x)=10x cents
Number of quarters is "45-x-2x" : Value of these is 25(45-3x)=1125-75x cents
----------
EQUATION:
Value + Value + Value = Value
x + 10x + 1125-75x=260 cents
-64x=260-1125
-64x = -865
Comment: This does not yield a whole number value for x.
Are you sure the account has $2.60 in it?
Cheers,
Stan H.